A 50 kg block sits on a level surface. One end of the surface is raised until the block begin to slide down the ramp at a constant velocity. What is the angle if the coefficient friction is 0.750 between the block and the incline plane in degrees?


1
Expert's answer
2021-03-18T14:01:51-0400

"mg\sin\alpha-\mu N=ma=0"


"N=mg\cos\alpha"


"mg\sin\alpha=\mu mg\cos\alpha \to \mu=\tan\alpha\to"


"\alpha=\tan^{-1}\mu=\tan^{-1}0.75\approx37°" . Answer